TC03E08 A chemical reaction carried out at 30 °C takes in 550 J of heat which is transferred from the surroundings. The...

TC03E08

A chemical reaction carried out at 30 °C takes in 550 J of heat which is transferred from the surroundings. The reaction produces a gaseous product so that the reaction does 280 J of work on the atmosphere as the chemicals expand. What is ΔU for this reaction?

Select one:

a. -830 J

b. +270 J

c. -270 J

d. -278 J

e. +830 J
Answer #1

apply first law of thermodynamics

ΔU = q + w

q= heat transferred , since the heat is transferred from the surrounding it means heat is absorbed by the system therefore it will be positive

q = + 550 J

w = work done on the atmosphere means work is done by the system . work done by system is negative

w = - 280 J

ΔU = + 550 + ( - 280) = 550 - 280 = 270 J

answer is option B +270 J